# Adam Exploits $\ell_{\infty}$-geometry of Loss Landscape via Coordinate-wise Adaptivity

### v1

@@ 3.3 A unified analysis for blockwise $\mathtt{Adam}$
Under 3.10, suppose $L$ is $\mathbf{H}$-smooth blockwisely w.r.t. $(\infty,2)$-norm, where $\mathbf{H}=(H_{1},\ldots,H_{B})\in\mathbb{R}^{B}$, for Algorithm 3, we have that
$\min_{\frac{T}{2}<t\leq T}\mathbb{E}\sum_{b=1}^{B}\sqrt{d_{b}}\left\|\nabla_{(b)}L({\bm{x}}_{t})\right\|_{2}$ $\leq E+\sqrt{E}\sqrt{\frac{\beta_{2}^{\frac{T}{4}}}{T(1-\beta_{2})}d\sqrt{v_{0}}+\sum_{b=1}^{B}d_{b}\sigma_{b}+d\sqrt{\epsilon}}$
@@
$E$ $=\frac{2}{\eta T}\mathbb{E}\left[L({\bm{x}}_{0})-L({\bm{x}}_{T})\right]+\left(1+\frac{\beta_{2}F}{T(1-\beta_{2})}\right)\left(\eta{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\sum_{b=1}^{B}H_{b}}+\sqrt{1-\beta_{2}}\sum_{b=1}^{B}d_{b}\sigma_{b}\right),$
@@
$F=2\ln\left(1+\frac{\sum_{b=1}^{B}\sigma_{b}^{2}+\left\|\nabla L({\bm{x}}_{0})\right\|_{\Phi}^{2}+\max_{b\in[B]}H_{b}^{2}\eta^{2}T(T+\frac{1}{1-\beta_{2}})}{v_{0}+\epsilon}\right)+\ln{32}.$
(1,1)-norm as a surrogate complexity measure for $H(L,\Phi_{\mathtt{Adam}})$.
$H_{i}$ in Definition 3.3 is determined by $\sup_{{\bm{x}}}\sum_{j=1}^{d}\left|\nabla^{2}_{i,j}L({\bm{x}})\right|$, which is difficult to compute because it requires taking supreme over the entire domain. An computationally-tractable alternative is to approximate $\sum_{i=1}^{d}H_{i}$ locally by the $(1,1)$-norm of Hessian of loss along the training trajectory. We provide an efficient approximation algorithm with guarantees by using hessian-vector product against random Cauchy vectors in Section D.2.
@@
As an implication of Theorem 3.11, we immediately get analogs of Corollaries 3.7, 3.7 and 3.6 for $\mathtt{AdaSGD}$, with the corresponding new noise assumption and smoothness assumption. When the optimization is not noise-dominated, i.e., the main term in the deterministic case $\sqrt{\frac{R}{T}}=\sqrt{\frac{(L({\bm{x}}_{0})-\min_{{\bm{x}}}L({\bm{x}}))H(L,\Phi)}{T}}$ becomes the leading term, the choice of $\Phi$ now matters a lot. Here the biggest change from $\mathtt{AdaSGD}$ to $\mathtt{Adam}$ is the difference between $H(L,\Phi_{\mathtt{AdaSGD}})$ and $H(L,\Phi_{\mathtt{Adam}})$, which roughly correspond to $\sup_{{\bm{x}}}\|\nabla^{2}L({\bm{x}})\|_{2}$ and $\sup_{{\bm{x}}}\|\nabla^{2}L({\bm{x}})\|_{1,1}$, using the above mentioned approximation.
Previous analyses of $\mathtt{Adam}$’s convergence (Shi & Li, 2021; Défossez et al., 2022; Li & Lin, 2024) usually assume smoothness under the $\ell_{2}$ norm. When using this assumption, the resulting convergence rate for $\mathtt{Adam}$ ends up being identical to the rate for $\mathtt{AdaSGD}$, which fails to capture why $\mathtt{Adam}$ often performs better than $\mathtt{AdaSGD}$ in practice. By shifting to an $\ell_{\infty}$ norm smoothness assumption, we can observe the key difference: the coefficient for $\mathtt{Adam}$’s convergence rate changes from $\sup_{{\bm{x}}}\|\nabla^{2}L({\bm{x}})\|_{2}$ to $\sup_{{\bm{x}}}\|\nabla^{2}L({\bm{x}})\|_{1,1}$, where the latter is typically much smaller when $\mathtt{Adam}$ optimizes faster. This change leads to a divergence in the rates of the two algorithms, and comparing these new coefficients can provide insight into which algorithm may be more effective under different conditions.
@@
We will use $\bar{{\bm{g}}}_{t}=\mathbb{E}[{\bm{g}}_{t}|{\bm{x}}_{<t}]=\nabla L({\bm{x}}_{t-1})$ to denote the full batch gradient. We first introduce Lemma 3.12, which is the main and only usage of our new smoothness assumption Definition 3.9.
Lemma 3.12.
For any twice differentiable loss which is $\mathbf{H}$-smooth block-wisely w.r.t. $(\infty,2)$-norm (Definition 3.9), we have for any ${\bm{x}}$ and ${\bm{\Delta}}\in\mathbb{R}^{d}$,
${\bm{\Delta}}^{\top}\nabla^{2}L({\bm{x}}){\bm{\Delta}}\leq\sum_{b=1}^{B}\frac{H_{b}}{d_{b}}\left\|{\bm{\Delta}}_{(b)}\right\|_{2}^{2}.$ (1)
We start by considering the decrease of $L({\bm{x}}_{t})$ in a single step $t$. We can upper bound the second order term in the Taylor expansion (Equation 2) with Lemma 3.12. Then we can get loss decrease in a single step
$L({\bm{x}}_{t})-L({\bm{x}}_{t-1})$ $\leq-\eta\sum_{i=1}^{d}\frac{g_{t,i}\bar{g}_{t,i}}{\sqrt{v_{t,\Phi(i)}+\epsilon}}+\frac{1}{2}\eta^{2}\sum_{b=1}^{B}\frac{H_{b}}{d_{b}}\frac{\left\|{\bm{g}}_{t,(b)}\right\|_{2}^{2}}{v_{t,b}+\epsilon}$ (2)
$\leq-\eta\sum_{b=1}^{B}\frac{{\bm{g}}_{t,(b)}^{\top}\bar{{\bm{g}}}_{t,(b)}}{\sqrt{v_{t,b}+\epsilon}}+\frac{1}{2}\eta^{2}\sum_{b=1}^{B}H_{b}\frac{\left\|{\bm{g}}_{t,(b)}\right\|_{2}^{2}/d_{b}}{v_{t,b}+\epsilon}$ (3)
The proof contains two main parts: lower bounding the first order term and upper bounding the second order term. Below we address the second-order term first. At each step, the second order term $\frac{\left\|{\bm{g}}_{t,(b)}\right\|_{2}^{2}/d_{b}}{v_{t,b}+\epsilon}$ can be as large as $\frac{1}{1-\beta_{2}}$. But if we sum over all the steps, we can employ Lemma 3.13 to bound the sum by $T+\frac{\beta_{2}}{1-\beta_{2}}\ln{\frac{v_{T,b}+\epsilon}{v_{0,b}+\epsilon}}$ rather than $\frac{T}{1-\beta_{2}}$.
@@
Now we turn to the first term. Ideally, we would like to connect the first order term to $\|\bar{\bm{g}}_{t,(b)}\|_{2}^{2}$ by taking expectation, e.g., $\mathbb{E}_{t}\frac{{\bm{g}}_{t,(b)}^{\top}\bar{{\bm{g}}}_{t,(b)}}{\sqrt{v_{t,b}+\epsilon}}\approx\frac{\mathbb{E}_{t}{\bm{g}}_{t,(b)}^{\top}\bar{{\bm{g}}}_{t,(b)}}{\sqrt{v_{t,b}+\epsilon}}=\frac{\|\bar{\bm{g}}_{t,(b)}\|_{2}^{2}}{\sqrt{v_{t,b}+\epsilon}}$, where we use $\mathbb{E}_{t}[\cdot]$ as abbreviation for $\mathbb{E}[\cdot|{\bm{x}}_{<t}]$. However, this is not correct because both the numerator and denominator in $\sum_{b=1}^{B}\frac{{\bm{g}}_{t,(b)}^{\top}\bar{{\bm{g}}}_{t,(b)}}{\sqrt{v_{t,b}+\epsilon}}$ depend on the stochastic gradient noise at step $t$. To circumvent this difficulty, we use Lemma 3.14 to lower bound each $\mathbb{E}_{t}\frac{{\bm{g}}_{t,(b)}^{\top}\bar{{\bm{g}}}_{t,(b)}}{\sqrt{v_{t,b}+\epsilon}}$ by $\frac{\mathbb{E}_{t}{\bm{g}}_{t,(b)}^{\top}\bar{{\bm{g}}}_{t,(b)}}{\sqrt{\mathbb{E}_{t}v_{t,b}+\epsilon}}\geq\frac{\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}}{\sqrt{\tilde{v}_{t,b}+\epsilon}}$ minus error terms related to noise magnitude $\sigma_{b}$, where $\tilde{v}_{t,b}=\beta_{2}v_{t-1,b}+(1-\beta_{2})\left(\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}/d_{b}+\sigma_{b}^{2}\right)$.
@@
For any block $b\in[B]$, it holds that
@@
Finally, we employ Cauchy inequality to upper bound $\left\|\bar{{\bm{g}}}_{t}\right\|_{\Phi,*}$:
@@
We use Lemma 3.15 to upper bound $\sum_{t=\frac{T}{2}+1}^{T}d_{b}\sqrt{\tilde{v}_{t,b}+\epsilon}$ by $\sum_{t=1}^{T}\frac{\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}}{\sqrt{\tilde{v}_{t,b}+\epsilon}}$. $\sum_{t=1}^{T}\sum_{b=1}^{B}\frac{\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}}{\sqrt{\tilde{v}_{t,b}+\epsilon}}$ can be bounded from analysis above, which finishes the proof.
@@
For any block $b\in[B]$, it holds that
@@ 4 Experiments
In order to empirically investigate and confirm the implications of our propsosed theory, we compare the training performance of $\mathtt{Adam}$ with $\mathtt{AdaSGD}$, $\mathtt{SGD}$ and $\mathtt{rotated\ Adam}$ on multiple different tasks. In $\mathtt{Adam}$ and its variants (including AdaSGD) we always set $\beta_{1},\beta_{2}=(0.9,0.99)$. Momentum in SGD is also set to $0.9$. Weight decay is always deactivated.
@@ 4.1 Quadratic loss
We perform controlled experiments on quadratic loss to study the relationship between optimization speed of $\mathtt{Adam}$ and the shape of Hessian, in terms of $\Phi_{\mathtt{Adam}}$-smoothness. More specifically, we consider $\Sigma=\text{diag}(\underbrace{1,\cdots,1}_{10},1,\frac{1}{2^{2}},\frac{1}{3^{2}},\cdots,\frac{1}{990^{2}})\in\mathbb{R}^{1000\times 1000}$ and optimize the corresponding quadratic loss of $\Sigma$ by $\mathtt{Adam}$, with different levels of rotations. We manually generate orthogonal matrices $\mathcal{R}_{i}$ in the following way. We first sample $\mathcal{M}\in\mathbb{R}^{d\times d}$ where $\mathcal{M}_{i,j}$ is i.i.d. sampled from $N(0,1)$. Then $\mathcal{A}=\mathcal{M}-\mathcal{M}^{\top}$ is a skew-symmetric matrix and $\exp{(t\mathcal{A})}$ represents a continuous family of matrices. We define $\mathcal{R}_{i}=\exp{(t_{i}\mathcal{A})}$ for different $t_{i}$. When $t_{i}=0$, we know $\mathcal{R}_{i}=I$. When $t_{i}\to\infty$, $\mathcal{R}_{i}$ converges to a random orthogonal matrix in distribution.
Then we optimize $L_{0}({\bm{x}})={\bm{x}}^{\top}\Sigma{\bm{x}}$ with $\mathtt{AdaSGD}$ and $\mathtt{Adam}$ and optimize $L_{i}({\bm{x}})={\bm{x}}^{\top}\mathcal{R}_{i}^{\top}\Sigma\mathcal{R}_{i}{\bm{x}}$ with $\mathtt{Adam}$ for $100$ steps. Because $\mathtt{AdaSGD}$ is rotational invariant, the optimization performance of $\mathtt{AdaSGD}$ is the same on all $L_{i}$. We tune learning rates for each setting and present their best results in Table 1.
We find a clear pattern that $\mathtt{Adam}$ optimizes worse when the $(1,1)$ norm of Hessian matrix increases, as suggested by our Corollary 3.7. Moreover, when $(1,1)$ norm divided by dimension is smaller than spectral norm, $\mathtt{Adam}$ tends to optimize faster than $\mathtt{AdaSGD}$ and vice versa, as suggested by our Theorem 3.11.
@@
AdaSGD $0.00582$ $0.00881$ $0.00172$
Adam $0.00582$ $0.00030$ $0.00001$
Adam ($\mathcal{R}_{1}$) $0.04162$ $0.00317$ $0.00062$
Adam ($\mathcal{R}_{2}$) $0.25364$ $0.00588$ $0.00122$
Adam ($\mathcal{R}_{3}$) $0.61866$ $0.00747$ $0.00179$
Adam ($\mathcal{R}_{4}$) $1.29959$ $0.00920$ $0.00239$
Table 1: The final loss values obtained by different optimizers and the $(1,1)$-norm of Hessian matrix for the corresponding unrotated objective and rotated objectives. The spectral norm of the Hessian matrix is always $1$. $\mathtt{Adam}$ optimizes worse when the $(1,1)$-norm of Hessian matrix increases, as suggested by our Corollary 3.7. Moreover, when $(1,1)$-norm is smaller than spectral norm times space dimension, $\mathtt{Adam}$ tends to optimize faster than $\mathtt{AdaSGD}$ and vice versa, which justifies the effectiveness of $\Phi$-smoothness as a tool to predict the optimization speed of blockwise $\mathtt{Adam}$.

### v3

@@ 3.3 A unified analysis for blockwise $\mathtt{Adam}$
For a specific partition $\Phi$, we consider the updates defined in Algorithm 3. Under 3.10, we have that
$\min_{\frac{T}{2}<t\leq T}\mathbb{E}\sum_{b=1}^{B}\sqrt{d_{b}}\left\|\nabla_{(b)}L({\bm{x}}_{t})\right\|_{2}$ $\leq 2\sqrt{2}E+\sqrt{2E}\sqrt{\frac{4\beta_{2}^{\frac{T}{4}}}{T(1-\beta_{2})}d\sqrt{v_{0}}+\sum_{b=1}^{B}d_{b}\sigma_{b}+d\sqrt{\epsilon}}$
@@
$E$ $=\frac{2}{\eta T}\mathbb{E}\left[L({\bm{x}}_{0})-L({\bm{x}}_{T})\right]+\left(1+\frac{\beta_{2}F}{T(1-\beta_{2})}\right)\left(\eta{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}H(L,\Phi)}+2\sqrt{1-\beta_{2}}\sum_{b=1}^{B}d_{b}\sigma_{b}\right),$
@@
$F=2\ln\left(1+\frac{\sum_{b=1}^{B}\sigma_{b}^{2}+\left\|\nabla L({\bm{x}}_{0})\right\|_{\Phi}^{2}+\sum_{b\in[B]}H_{b}^{2}d_{b}\eta^{2}T(T+\frac{1}{1-\beta_{2}})}{v_{0}+\epsilon}\right)+\ln{32}.$
In order to obtain Theorem 3.5 from Theorem 3.11, we rely on the following definition that generalizes Definition 3.3.
Definition 3.12 (Generalized version of Definition 3.3).
For any partition function $\Phi:[d]\rightarrow[B]$ and ${\bm{H}}=(H_{1},\ldots,H_{B})\in\mathbb{R}^{B}$, we say a function $L$ is $\mathbf{H}$-blockwisely-smooth w.r.t. $\Phi$-norm , iff for any $b\in[B]$, ${\bm{x}},{\bm{y}}\in\mathbb{R}^{d}$,
$\left\|\nabla_{(b)}L({\bm{x}})-\nabla_{(b)}L({\bm{y}})\right\|_{2}\leq\sqrt{d_{b}}H_{b}\left\|{\bm{x}}-{\bm{y}}\right\|_{\Phi}$
The following Lemma 3.13 will show that $H(L,\Phi)$ is upper bounded by $\sum_{b=1}^{B}d_{b}H_{b}$ when $L$ is ${\bm{H}}$-blockwisely-smooth w.r.t. $\Phi$-norm. The proof of Lemma 3.13 is deferred in Appendix C.
Lemma 3.13.
For any twice differentiable loss which is $\mathbf{H}$-blockwisely-smooth w.r.t. $\Phi$-norm (Definition 3.12), we have for any ${\bm{x}}$ and ${\bm{\Delta}}\in\mathbb{R}^{d}$,
$\left|{\bm{\Delta}}^{\top}\nabla^{2}L({\bm{x}}){\bm{\Delta}}\right|\leq\sum_{b=1}^{B}H_{b}\left\|{\bm{\Delta}}_{(b)}\right\|_{2}^{2}.$ (2)
Then the diagonal matrix ${\bm{A}}$ defined by ${\bm{A}}_{i,i}=H_{\Phi(i)}$ follows partition $\Phi$ and always dominates $\left|\nabla^{2}L({\bm{x}})\right|$.
When $L$ is $(H_{1},\dots,H_{d})$-smooth coordinate-wisely w.r.t. $\ell_{\infty}$ norm, it satisfies ${\rm diag}(H_{1},\dots,H_{d})\succeq\left|\nabla^{2}L({\bm{x}})\right|$ for all ${\bm{x}}\in\mathbb{R}^{d}$ and ${\rm diag}(H_{1},\dots,H_{d})$ follows $\Phi_{\mathtt{Adam}}$ partition. So $H(L,\Phi_{\mathtt{Adam}})$ is at most $\sum_{i=1}^{d}H_{i}$ and we can directly derive Theorem 3.5 from Theorem 3.11. For $\Phi_{\mathtt{AdaSGD}}$ being the mapping $i\mapsto 1$, $H(L,\Phi_{\mathtt{AdaSGD}})$ is the same as the smoothness under $\Phi_{\mathtt{AdaSGD}}$-norm, whose value is $d\sup_{{\bm{x}}\in\mathbb{R}^{d}}\left\|\nabla^{2}L({\bm{x}})\right\|_{2}$.
@@
As an implication of Theorem 3.11, we can get analogs of Corollaries 3.7, 3.7 and 3.6 for $\mathtt{AdaSGD}$, with the corresponding noise and smoothness assumptions. When the optimization is not noise-dominated and $\sqrt{\frac{R}{T}}=\sqrt{\frac{(L({\bm{x}}_{0})-\min_{{\bm{x}}}L({\bm{x}}))H(L,\Phi)}{T}}$ becomes the leading term, the choice of $\Phi$ now matters a lot. The key difference between $\mathtt{AdaSGD}$ and $\mathtt{Adam}$ lies in the gap between $H(L,\Phi_{\mathtt{AdaSGD}})$ and $H(L,\Phi_{\mathtt{Adam}})$ and comparing these coefficients can provide insight into which algorithm may be more effective under different conditions.
Previous analyses of $\mathtt{Adam}$’s convergence (Shi & Li, 2021; Défossez et al., 2022; Li & Lin, 2024) usually assume smoothness under the $\ell_{2}$ norm and the rate of $\mathtt{Adam}$ ends up being identical to the rate of $\mathtt{AdaSGD}$, which fails to explain why $\mathtt{Adam}$ often performs better than $\mathtt{AdaSGD}$ in practice. By adopting an $\ell_{\infty}$ norm smoothness assumption, the coefficient for $\mathtt{Adam}$’s convergence rate changes from $d\sup_{{\bm{x}}}\|\nabla^{2}L({\bm{x}})\|_{2}$ to $H(L,\Phi_{\mathtt{Adam}})$, where the latter is typically much smaller when $\mathtt{Adam}$ optimizes faster because $\sup_{\bm{x}}\left\|\nabla^{2}L({\bm{x}})\right\|_{1,1}$ is much smaller than $d\sup_{\bm{x}}\left\|\nabla^{2}L({\bm{x}})\right\|_{2}$.
@@
We will use $\bar{{\bm{g}}}_{t}=\mathbb{E}[{\bm{g}}_{t}|{\bm{x}}_{<t}]=\nabla L({\bm{x}}_{t-1})$ to denote the full batch gradient. We start by considering the decrease of $L({\bm{x}}_{t})$ in a single step $t$. We can upper bound the second order term in the Taylor expansion (Equation 3) with ${\bm{H}}\succeq\nabla^{2}L({\bm{x}})$ that achieves $\operatorname{Tr}({\bm{H}})=H(L,\Phi)$. Then we can get
$L({\bm{x}}_{t})-L({\bm{x}}_{t-1})$ $\leq-\eta\sum_{i=1}^{d}\frac{g_{t,i}\bar{g}_{t,i}}{\sqrt{v_{t,\Phi(i)}+\epsilon}}+\frac{1}{2}\eta^{2}\sum_{b=1}^{B}H_{b}\frac{\left\|{\bm{g}}_{t,(b)}\right\|_{2}^{2}}{v_{t,b}+\epsilon}$ (3)
$\leq-\eta\sum_{b=1}^{B}\frac{{\bm{g}}_{t,(b)}^{\top}\bar{{\bm{g}}}_{t,(b)}}{\sqrt{v_{t,b}+\epsilon}}+\frac{1}{2}\eta^{2}\sum_{b=1}^{B}H_{b}d_{b}\frac{\left\|{\bm{g}}_{t,(b)}\right\|_{2}^{2}/d_{b}}{v_{t,b}+\epsilon}$ (4)
The proof contains two main parts: lower bounding the first order term using $\|\bar{{\bm{g}}}_{t}\|_{\Phi,*}$ and upper bounding the second order term. Below we address the second-order term first. As mentioned in Section 3.2, the second order term $\frac{\left\|{\bm{g}}_{t,(b)}\right\|_{2}^{2}/d_{b}}{v_{t,b}+\epsilon}$ can be as large as $\frac{1}{1-\beta_{2}}$ for one step. But we can employ Lemma 3.14 to bound the sum by $T+\frac{\beta_{2}}{1-\beta_{2}}\ln{\frac{v_{T,b}+\epsilon}{v_{0,b}+\epsilon}}$ rather than $\frac{T}{1-\beta_{2}}$, where we set $v_{t}\triangleq v_{t,b}$ and $g_{t}\triangleq\left\|{\bm{g}}_{t,(b)}\right\|_{2}/\sqrt{d_{b}}$.
@@
Now we turn to the first term. Ideally, for each block $b\in[B]$, we would like to connect the first order term to $\|\bar{\bm{g}}_{t,(b)}\|_{2}^{2}$ by taking expectation, i.e., $\mathbb{E}_{t}\frac{{\bm{g}}_{t,(b)}^{\top}\bar{{\bm{g}}}_{t,(b)}}{\sqrt{v_{t,b}+\epsilon}}\approx\frac{\mathbb{E}_{t}{\bm{g}}_{t,(b)}^{\top}\bar{{\bm{g}}}_{t,(b)}}{\sqrt{v_{t,b}+\epsilon}}=\frac{\|\bar{\bm{g}}_{t,(b)}\|_{2}^{2}}{\sqrt{v_{t,b}+\epsilon}}$, where we use $\mathbb{E}_{t}[\cdot]$ as abbreviation for $\mathbb{E}[\cdot|{\bm{x}}_{<t}]$. However, this is not correct because both the numerator and denominator in $\frac{{\bm{g}}_{t,(b)}^{\top}\bar{{\bm{g}}}_{t,(b)}}{\sqrt{v_{t,b}+\epsilon}}$ depend on the stochastic gradient ${\bm{g}}_{t}$. To circumvent this difficulty, we lower bound each conditional expectation $\mathbb{E}_{t}\frac{{\bm{g}}_{t,(b)}^{\top}\bar{{\bm{g}}}_{t,(b)}}{\sqrt{v_{t,b}+\epsilon}}$ by $\frac{\mathbb{E}_{t}{\bm{g}}_{t,(b)}^{\top}\bar{{\bm{g}}}_{t,(b)}}{2\sqrt{\mathbb{E}_{t}v_{t,b}+\epsilon}}$, minus error terms related to noise magnitude $\sigma_{b}$. We can further have $\frac{\mathbb{E}_{t}{\bm{g}}_{t,(b)}^{\top}\bar{{\bm{g}}}_{t,(b)}}{\sqrt{\mathbb{E}_{t}v_{t,b}+\epsilon}}\geq\frac{\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}}{\sqrt{\tilde{v}_{t,b}+\epsilon}}$, where $\tilde{v}_{t,b}=\beta_{2}v_{t-1,b}+(1-\beta_{2})\left(\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}/d_{b}+\sigma_{b}^{2}\right)$. This leads to Lemma 3.15 whose proof is deferred to Appendix C.
@@
With 3.10, it holds that for any block $b\in[B]$
@@
Combining Lemmas 3.14 and 3.15 and Equation 4 yields an upper bound for $\mathbb{E}\sum_{t=1}^{T}\sum_{b=1}^{B}\frac{\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}}{\sqrt{\tilde{v}_{t,b}+\epsilon}}$ involving initial suboptimality gap, noise and smoothness. Finally, we employ Cauchy inequality (Equation 7) and Lemma 3.16 to upper bound $\left\|\bar{{\bm{g}}}_{t}\right\|_{\Phi,*}$ using $\mathbb{E}\sum_{t=\frac{T}{2}+1}^{T}\sum_{b=1}^{B}\frac{\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}}{\sqrt{\tilde{v}_{t,b}+\epsilon}}$. This completes the proof.
@@
With 3.10, it holds that for any block $b\in[B]$
@@ 4 Experiments
In order to empirically investigate and confirm the implications of our proposed theory, we compare the training performance of $\mathtt{Adam}$ with $\mathtt{AdaSGD}$, $\mathtt{SGD}$ and $\mathtt{rotated\ Adam}$ on multiple different tasks. In $\mathtt{Adam}$ and its variants (including AdaSGD) we set $(\beta_{1},\beta_{2})=(0.9,0.99)$ for experiments on the quadratic loss (Section 4.1) and ResNet18 (Section 4.3) and $(\beta_{1},\beta_{2})=(0.9,0.95)$ for experiments on GPT-2 (Section 4.2). Momentum in SGD is also set to $0.9$. Weight decay is always deactivated.
@@ 4.1 Quadratic loss
We perform controlled experiments on quadratic loss to study the relationship between optimization speed of $\mathtt{Adam}$ and the shape of Hessian in terms of $\Phi_{\mathtt{Adam}}$-smoothness. More specifically, we consider $\Sigma=\text{diag}(\underbrace{1,\cdots,1}_{10},1,\frac{1}{2^{2}},\frac{1}{3^{2}},\cdots,\frac{1}{990^{2}})\in\mathbb{R}^{1000\times 1000}$ and manually generate orthogonal matrices $\mathcal{R}_{i}$ in the following way. We first sample $\mathcal{M}\in\mathbb{R}^{d\times d}$ where $\mathcal{M}_{i,j}$ is i.i.d. sampled from $N(0,1)$. Then $\mathcal{A}=\mathcal{M}-\mathcal{M}^{\top}$ is a skew-symmetric matrix and $\exp{(t\mathcal{A})}$ represents a continuous family of matrices. We define $\mathcal{R}_{i}=\exp{(t_{i}\mathcal{A})}$ for different $t_{i}$. When $t_{i}=0$, we know $\mathcal{R}_{i}=I$. When $t_{i}\to\infty$, $\mathcal{R}_{i}$ converges to a random orthogonal matrix in distribution. We pick $t_{1}=0.002,t_{2}=0.008,t_{3}=0.015,t_{4}=0.1$ for our experiments.
Then we optimize $L_{0}({\bm{x}})=\frac{1}{2}{\bm{x}}^{\top}\Sigma{\bm{x}}$ with $\mathtt{AdaSGD}$ and $\mathtt{Adam}$ and optimize $L_{i}({\bm{x}})=\frac{1}{2}{\bm{x}}^{\top}\mathcal{R}_{i}^{\top}\Sigma\mathcal{R}_{i}{\bm{x}}$ with $\mathtt{Adam}$ for $100$ steps. Because $\mathtt{AdaSGD}$ is rotation-equivariant, the optimization process of $\mathtt{AdaSGD}$ is the same on all $L_{i}$. The initial ${\bm{x}}_{0}$ is decided by sampling from $\text{Unif}([0,1]^{1000})$ when there is no rotation. When $\mathcal{R}_{i}$ is not identity matrix, we will start training from $\mathcal{R}_{i}^{\top}{\bm{x}}_{0}$ to ensure the initial loss values are the same across different rotations. We tune learning rates for each setting with $10$ random seeds and present their lowest average loss with standard deviation in Table 1.
We find a clear pattern that $\mathtt{Adam}$ optimizes worse when the $(1,1)$-norm of Hessian matrix increases, as suggested by our Corollary 3.7. Moreover, when $(1,1)$-norm divided by dimension is smaller than spectral norm, $\mathtt{Adam}$ tends to optimize faster than $\mathtt{AdaSGD}$, as suggested by our Theorem 3.11.
@@
AdaSGD $0.01164$ $0.00887\pm 0.00119$ $0.00405\pm 0.00021$
Adam $0.01164$ $0.00022\pm 0.00007$ $0.00002\pm 0.00001$
Adam ($\mathcal{R}_{1}$) $0.08324$ $0.00314\pm 0.00031$ $0.00066\pm 0.00008$
Adam ($\mathcal{R}_{2}$) $0.50729$ $0.00567\pm 0.00053$ $0.00134\pm 0.00007$
Adam ($\mathcal{R}_{3}$) $1.23731$ $0.00751\pm 0.00086$ $0.00183\pm 0.00009$
Adam ($\mathcal{R}_{4}$) $2.59919$ $0.00978\pm 0.00132$ $0.00254\pm 0.00008$
Table 1: The final loss values obtained by different optimizers and the $(1,1)$-norm of Hessian matrix for the corresponding unrotated objective and rotated objectives. The spectral norm of the Hessian matrix is always $1$. $\mathtt{Adam}$ optimizes worse when the $(1,1)$-norm of Hessian matrix increases, as suggested by our Corollary 3.7. Moreover, when $(1,1)$-norm divided by $d$ is much smaller than spectral norm, $\mathtt{Adam}$ tends to optimize faster than $\mathtt{AdaSGD}$, which justifies the effectiveness of $\Phi$-smoothness as a tool to predict the optimization speed of blockwise $\mathtt{Adam}$.